OpenStudy (anonymous):

A jet plane traveling at 550 mph overtakes a propeller plane traveling at 200 mph that had a 2-hour head start. How far from the starting point are the planes?

OpenStudy (anonymous):

Equation: distance = distance 500(x-2) = 200x 500x - 1000 = 200x 300x = 1000 x = 3 1/3 hrs (time the propeller plane flew before being overtaken) Propeller distance = 200x = 200*(10/3) = 666.67 miles